INAF LBC+LUCIFER queue observing, 2010 December 01 - 02

Observer: K.Boutsia, M. Fumana,V. Testa
Telescope Operator: S. Allanson
Support Astronomer: D. Thompson
Instrument: LBC - LUCIFER

Summary

Thin clouds but generally good conditions.
Seeing <0.8''
Attention: Flats have been named luci_20101011_Flats and are difficult to find in archive. We saved locally and renamed them.

Data Summary

Project Objects Filter ExpTime Notes
Clementini AndXXI - AndXIX B-bessel, V-bessel 105min + 105min seeing <1.0' - some elongation problems at the beginning
Turatto SN2009kn J, H, Ks 640 s + 896 s + 896 s ??????
Turatto SN2010jl J, H, Ks 12 s + 12 s + 24 s repeated with different guide star
Turatto SN2010gi J, H, Ks 80 s + 160 s +240 s low airmass
Cresci HDF 210zJHK+H 1hr ARC+FLAT - Tell
Cresci HDF 210zJHK+K 1hr ARC+FLAT - Tell

Details

Begin night with LBC

12:25 Turn on LBC
  • 10Bias_Bino

12:34 - Skyflat Test LBC - B(blue) - V(red) ok

12:44 - SkyFlat B-V
  • PA 180 2.5 sec B - 2.5sec V
  • PA 90 - 6sec B + 6sec V
  • PA 270 10sec B - 10sec V

12:53 - Slew to pointing star WT10_351
  • Pointing correction OK + mirror adjust OK
  • problems with setting the camera name - David fixed it again
  • it does not read LBC or RedLBC as lbcr and it does not find the file

01:16 - slew to Clementini - AndXXI -dofpia
  • manually give correction becuase of large come in the Red and some astigmatism in the Blue
  • hit limits on red mirror

01:39 - observing target Clementini - M31_AndXXI_420s_6
  • Seeing good - 0.8'' from the dimm
  • Ob aborted because focus was lost on the red mirror. *dofpia

01:56 - repeat OB - M31_AndXXI_420s_6
  • seeing 0.8'' on both sides
  • blue images slightly elongated

02:20 - slew to AndXIX - dofpia

02:26 - observing target Clementini - M31_AndXIX_420s_6
  • very high elevation - it can not keep focus

02:50 - slew to AndXXI - dofpia

03:00 - observing target Clementini - M31_AndXXI_420s_7

03:24 - slew to AndXIX - dofpia

03:31 - observing target Clementini - M31_AndXIX_420s_7

03:56 - Slew to AndXXI - dofpia
  • after 5 iterations still not converging with 0.7 seeing.....
  • stopped with small residual X astig on SX

04:08 - observing target Clementini - M31_AndXXI_420s_8
  • tonight left mirror is having big trouble in keeping itself collimated

04:31 - slew to AndXIX - dofpia
  • since the seeing is very good, we change the limits for dofpia (Initialize)
  • we have also put higher limits for the astigmatism since it doesn't seem to be able to correct it
  • dofpia, RedInitialize=250,BlueInitialize=300,Redlimit=300,Bluelimit=300

04:40 - observing target Clementini - M31_AndXIX_420s_8

05:05 - slew to AndXXI - dofpia
  • the rotator is very slow
  • 3 iterations but very good Zernikes

05:12 - observing target Clementini - M31_AndXXI_420s_9

05:38 - slew to AndXIX - dofpia
  • the rotator is very slow again
  • TO adjusted the mirrors
  • 3 iterations but very good Zernikes, bluie only two, expected better from the red
  • sky look ok, probably photometric

05:43 - observing target Clementini - M31_AndXIX_420s_9

06:08 - slew to AndXXI - dofpia
  • always with dofpia, RedInitial=250,BlueInitial=300,RedLimit=300,BlueLimit=300
  • blue converged during the 1st iteration - 2 iterations for the red

06:13 - observing target Clementini - M31_AndXXI_420s_10 * ended at airmass=1.51

06:36 - slew to AndXIX - dofpia
  • always with dofpia, RedInitial=250,BlueInitial=300,RedLimit=300,BlueLimit=300
  • RED converged during the 1st iteration - 2 iterations for the BLUE

06:41 - observing target Clementini - M31_AndXIX_420s_10
  • ended at airmass 1.58, seeing around 0.8 both filters

07:04 - Do standard SA92_BV (airmass~2.0)

07:09 - Do standard SA95_BV (airmass~1.2)

07:14 - Do standard SA98_BV (airmass~1.4)

07:18 - SWITCH TO LUCIFER

07:42 - Slew to SN2010gi for pointing correction and collimation

07:46 - acquisition imaging SN2010gi 0001-0002

07:53 - Imaging Science SN2010gi
  • H 0003-0006
  • J 0007-0010
  • Ks 0011-0014

08:09 - repeat imaging on SN2010jl: acquisition 0015-0016

08:16 - Imaging Science SN2010jl
  • H 17-0020
  • J 0021-0024
  • Ks 0025-0028

08:25 - Slew to SN2009kn Image Acquisition 0029-0030

08:29 - Imaging Science SN2009kn
  • H 0031-0046
  • J 0047-0054
  • Ks 0055-0070

09:35 - Acquisition Cresci HDF-N MOS
  • needed to change guide star, 10 minutes lost
  • mask turnout 0071-0072
  • through-slit 0073-0074

10:03 - MOS Science HDFN H_2 0075-0086

11:10 - Calibration Unit IN
  • ARC H 0087-0090
  • FLAT H 0091-0096

11:20 - Slew to Tell for HDF
  • repeat acquisition because guide star had to be changed
  • 0097-0103

11:34 - Sci Int for Tell in H (HDF)
  • too bright
  • the read mode is mer so even if DIT is 2sec it actually exposes for 10sec 0104-0111
  • image of Tell in the left most position without the grism to see if it is well centered 0112-0113

11:43 - Mirror adjustment and pointing correction

11:52 - Tell acq again for K band 0114-0115

11:57 - Sci Int for Tell in K (HDF) 0116-0123 * Telluric has no archive info in the header - just INAF

12:00 - Acquisition script for MOS spectrooscopy for Cresci (HDF) 0124-0128

12:15 - MOS science script for HDF in K band 0129-0140

13:21 - Calibration Unit IN
  • ARC K 0141-0144
  • FLAT K with Halo1 0145-0152
  • FLAT K with Halo2 0153-0158

13:34 - Calib Unit OFF - moving to flat position
  • Flats in J: 001-010 low (5000counts), 021-030 high (15000counts)
  • Flats in H:==011-020== low (5000counts), 031-040 high (15000-20000counts)
  • Flats in K: 041-070 (4000 - 8000counts)

13:53 - CLOSE DOME
  • darks script

-- VincenzoTesta - 02 Dec 2010
Topic revision: r7 - 15 Mar 2012, VincenzoTesta
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback